A stone is thrown upward and reaches a height of 19.6 meters. With what initial velocity was it thrown?

1
Expert's answer
2021-12-16T11:32:07-0500

"mgh=\frac{mV^2}{2}"

"V=\sqrt{2gh}=\sqrt{2\times 9.8 \times 19.6}=19.6m/s"